A population of squirrels may be brown (the dominant phenotype) or red (the recessive phenotype). Brown squirrels have the genotype BB or Bb. Red squirrels have the genotype bb. The frequency of the BB genotype is 0.46.


A. What is the frequency of heterozygous rabbits?


B. What is the frequency of the B allele? 


C. What is the frequency of the b allele?



1
Expert's answer
2021-02-08T10:21:32-0500

The frequencies of brown and red alleles are:

f (B) = f(BB)1/2 = 0.461/2 = 0.69

f (b) = 1 - f(B) = 1 - 0.69 = 0.31

A. The frequency of heterozygous squirrels are:

f (Bb) = 2 × f(B) × f(b) = 2 × 0.69 × 0.31 = 0.43

B. The frequency of the B allele:

f (B) = f(BB)1/2 = 0.461/2 = 0.69

C. The frequency of the b allele:

f (b) = 1 - f(B) = 1 - 0.69 = 0.31